OpenStudy (anonymous):

In triangle ABC shown below, use an appropriate trigonometric ratio to find the measurement of angle B. Round to the nearer degree.

OpenStudy (anonymous):



OpenStudy (anonymous):

It's B

OpenStudy (anonymous):

use cos

OpenStudy (anonymous):

\[\cos \theta=4/5\]

OpenStudy (anonymous):

now find theta

OpenStudy (anonymous):

hint: use cos^-1 on both sides of the equation

OpenStudy (anonymous):

@10115658 I'm afraid my calculator isn't working and I've been having difficulty understanding trigonometry...if you're not busy, could you walk me through the steps, please? Thank you so much in advance and I'm sorry for the inconvenience!

OpenStudy (anonymous):

what are you confused about?

OpenStudy (anonymous):

I don't understand how to find the measurements of angles using trigonometric ratios. I like math, and I don't mind studying it, but when I do study it it takes me a long time for things to click in. It's not my weakest subject, but it only becomes one of my stronger suits once I understand the formulas and whatnot of whatever section of math I'm learning. (In this case, trigonometry.) (^_^;)

OpenStudy (anonymous):

ok so the most important thing to remember is when to use sin, cos and tan

OpenStudy (anonymous):

use sin when you have the opposite side of a triangle and the hypotenuse

OpenStudy (anonymous):

SOH-CAH-TOA, right?

OpenStudy (anonymous):